# Disentangling representations of retinal images with generative models

### v1

@@ 3.3 Disentanglement loss
If the predictive model (Fig. 3) was only composed of an encoder and subspace classifiers, subspaces could correlate and share information. Therefore, we additionally penalized the presence of shared information with a disentanglement loss by minimizing the distance correlation (dCor) between subspaces.
Distance correlation (Székely et al., 2007) measures the dependence between random vectors. It is analogous to Pearson’s correlation but can also measure nonlinear dependencies between random vectors of arbitrary dimension. Moreover, the distance correlation is zero only if the random vectors are independent. We provide a toy example with linear and nonlinear dependencies in Appendix 6.1, where we compared the performance of minimizing distance correlation with minimizing linear dependence measures. In contrast to mutual information, distance correlation is easy to compute and does not require bound estimators. Furthermore, compared to adversarial classifiers, distance correlation showed better scalability when considering multiple subspaces, as adversarial classifiers add a learning problem for each subspace. For a more detailed background on dependence estimators we refer to the related work Sec. 2.
In practice, we computed the sample distance correlation measure (Székely et al., 2007) to disentangle subspaces with batch samples:
$\text{dCor}(w_{1},w_{2})$ $=\sqrt{\frac{\text{dCov}^{2}(w_{1},w_{2})}{\left(\text{dVar}^{2}(w_{1})\,\text{dVar}^{2}(w_{2})\right)^{1/2}}}$ (3)
We assumed n-dimensional batches of subspace vectors $w_{1}\in\mathbb{R}^{n\times d_{1}}$ and $w_{2}\in\mathbb{R}^{n\times d_{2}}$. We first mapped the subspace vectors into an $n\times n$ dimensional Euclidean space containing all pairwise distances from a batch:
$A\in\mathbb{R}^{n\times n},\,a_{j,k}$ $=\|w_{1}^{(j)}-w_{1}^{(k)}\|_{2},\,j,k=1,2,\dots,n$ (4)
$B\in\mathbb{R}^{n\times n},\,b_{j,k}$ $=\|w_{2}^{(j)}-w_{2}^{(j)}\|_{2},\,j,k=1,2,\dots,n$ (5)
The distance matrices for each subspace were then individually normalized by subtracting the row mean, column mean, and grand mean:
$A_{j,k}$ $\coloneqq a_{j,k}-\bar{a}_{j,\bullet}-\bar{a}_{\bullet,k}-\bar{a}_{\bullet,\bullet}$ (7)
$B_{j,k}$ $\coloneqq b_{j,k}-\bar{b}_{j,\bullet}-\bar{b}_{\bullet,k}-\bar{b}_{\bullet,\bullet}$ (8)
Based on the normalized subspace distance matrices, the sample squared distance covariance was simply the arithmetic average of the element-wise matrix product
$\text{dCov}^{2}(w_{1},w_{2})$ $=\frac{1}{n^{2}}\sum_{j=1}^{n}\sum_{k=1}^{n}A_{j,k}B_{j,k}$ (9)
and the sample squared distance variance was the arithmetic average of the squared matrix elements
$\text{dVar}^{2}(w_{1})$ $=\frac{1}{n^{2}}\sum_{j,k}A_{j,k}^{2}$ (10)
$\text{dVar}^{2}(w_{2})$ $=\frac{1}{n^{2}}\sum_{j,k}B_{j,k}^{2}$ (11)
The distance correlation of two random variables was finally obtained by dividing the distance covariance by the product of the distance standard deviations (Eq. 3).
Hence, we extended the optimization problem to learn disentangled subspaces with an encoder architecture as follows:
@@
$L_{\text{DC}}(\theta)$ $=\frac{1}{K}\sum_{i\neq k}^{K}\text{dCor}(w_{k},w_{i})$ (13)
3.4 Measuring disentanglement
Measuring disentanglement in a latent space is a non-trivial task. While there is no universally accepted definition of disentanglement, but most agree on two main points (Carbonneau et al., 2022). Firstly, there is a consensus that representation factors should exhibit independence, a quality commonly denoted as modularity. This implies that a factor affects only a subset of the representation space, and only this factor affects this subspace. Moreover, the affected representation space by a factor should be as small as possible, a principle known as compactness. The last agreed-upon criterion is the information content or explicitness. In essence, a disentangled representation should completely describe all factors of interest. To assess disentanglement, three distinct families of metrics have been proposed: intervention-based, predictor-based, and information-based metrics (Carbonneau et al., 2022).
In our case, we opted for a predictor-based metric to evaluate both modularity and explicitness, excluding compactness from consideration because we fixed subspace dimensions. For our predictor, we employed a k-nearest neighbors (kNN) classifier with k=30. Using this classifier, we generated a confusion matrix, comparing subspaces with associated labels, for example ethnicity and camera (Fig. 4). Given the presence of unbalanced classes, we reported the accuracy improvement over chance level accuracy. With the confusion matrix, modularity is evident through high classification accuracy improvement for one factor (e.g. ethnicity) in only one subspace. Explicitness, on the other hand, becomes apparent when the accuracy improvement is notably high across the entire matrix. Consequently, when both modularity and explicitness are effectively addressed, the confusion matrix has high values along the main diagonal as well as low values on the off-diagonals (Fig. 4).
@@ 3.5 Generative image model
For the generative model, the overall goal was to generate realistic fundus images with potentially correlated labels from a disentangled latent space. We considered the image generation task by learning the mapping between two domains $M:\mathcal{X}\rightarrow\mathcal{\hat{X}}$ with a latent space model. We split $M$ into two components, $M=d\circ g$, where the encoder/discriminator $d$ maps to an intermediate latent representation $d:\mathcal{X}\rightarrow\mathcal{W}$ and the generator $g:\mathcal{W}\rightarrow\mathcal{\hat{X}}$ maps to the output image domain.
@@
We chose to work with a StyleGAN2 architecture (Karras et al., 2020), a state-of-the-art generative adversarial network for high-fidelity and high-resolution images. The StyleGAN architecture has two features that makes it a potentially good candidate for disentanglement: (1) it consists of a mapping network $m:\mathcal{Z}\rightarrow\mathcal{W}$ that maps a latent vector into an intermediate space, which then controls the styles in each convolutional layer in the generator with adaptive instance normalization, and (2) it allows for the separation of fine-grained and coarse-grained features by a progressively growing generator.
@@
We trained the generative image model with the standard minimax GAN loss together with StyleGAN’s path length regularization (Karras et al., 2020) on the generator and an R1 regularizer (Mescheder et al., 2018) as a gradient penalty on the discriminator for real data. The resulting optimization problem can then be expressed as:
$L_{\text{GAN}}(\eta,\tau)$ $=\,\mathbb{E}_{x\in\mathbb{X}}\left[\log(d_{\tau,0}(x))\right]+\mathbb{E}_{z\in\mathbb{Z}}\left[\log(1-d_{\tau,0}(g_{\eta}(w)))\right]+\lambda_{PL}PL(\eta)-\lambda_{R_{1}}R_{1}(\tau)$ (14)
$\eta^{*}$ $=\operatorname*{arg\,min}_{\eta}\,L_{\text{GAN}}(\eta,\tau^{*})$ (15)
$\tau^{*}$ $=\operatorname*{arg\,min}_{\tau}\,-L_{\text{GAN}}(\eta^{*},\tau)$ (16)
@@
Here, $d_{\tau,0}(x)$ is the discriminator’s estimate of the probability that the real data instance $x$ is real, $g_{\eta}(w)$ denotes the generator’s output given $w$, and $d_{\tau,0}(g_{\eta}(w))$ is the discriminator’s estimate of the probability that a fake instance is real (Fig. 5).
@@
Figure 5: Generative image model with subspace constraints. For each image, we sample three individual latent codes from a standard normal distribution $z_{k}\sim\mathcal{N}(0,I_{d_{k}})$ and employ individual mapping networks $m_{\eta}^{k}:\mathcal{Z}\rightarrow\mathcal{W}_{k}$ to map them to intermediate latent spaces. The concatenation of all three subspaces $w=[w_{1},w_{2},w_{3}]\in\mathbb{R}^{d_{1}+d_{2}+d_{3}}$ serves as input to the generator $g_{\eta}:\mathcal{W}\rightarrow\mathcal{X}$. For GAN-inversion, we extend our discriminator with two additional heads: one for latent representation $\hat{w}$ and another for a pixel feature vector $p$ responsible for image encoding and pixel space reconstruction, respectively. For subspace learning, we introduce an individual mapping network for each subspace and incorporate label information for real images through classification heads for the first two subspaces. To prevent subspace correlation, we implement a penalty for the classification tasks using the average distance correlation (dCor) between the subspaces $\hat{w}_{k}$ that the discriminator outputs.
@@
As GANs were not originally designed for representation learning, they do not provide an encoder model. Therefore, we extended the StyleGAN2 architecture with an encoder to embed real images in the latest space for representation learning. Inspired by the work of invGAN (Ghosh et al., 2022), we extended the discriminator and also trained it as an encoder. We added two independent, fully connected layers and returned an estimate of the latent representation $\hat{w}$ for our first inversion loss
$L_{w}(\eta,\tau)$ $=\mathbb{E}_{z\in\mathbb{Z}}\left[||w-\hat{w}||_{2}^{2}\right]=\mathbb{E}_{z\in\mathbb{Z}}\left[\|w-d_{\tau,1}(g_{\eta}(w))\|_{2}^{2}\right].$ (17)
@@
For subspace learning, we modified the StyleGAN2 architecture to sample individual latent codes from a standard normal distribution $z_{k}\sim\mathcal{N}(0,I_{d_{k}})$ and learned individual 8-layer mapping networks for each subspace $m_{\eta}^{k}:\mathcal{Z}_{k}\rightarrow\mathcal{W}_{k}$ (Fig. 5, left), where each subspace can have a different dimensionality $w_{k}\in\mathbb{R}^{d_{k}}$. In setting up the generative image model, we assumed that we do not know all the underlying factors of variation or have access to their labels. Therefore, we provided the model with a free patient-identity subspace $w_{3}$ (Fig. 5), in which the encoder/discriminator can capture for further information required to reconstruct high-resolution retinal fundus images. As in the encoder model (Sec. 3.2), we encoded label information for real images by training linear classification heads $C_{\psi_{k}}$ for the first two subspaces (Fig. 5, right). To avoid subspace correlation, we penalized the classification task with the mean of all distance correlation measures between subspaces $\hat{w}_{k}$ that the discriminator outputs.
@@
$(\tau^{*},\psi^{*})$ $=\operatorname*{arg\,min}_{\tau,\psi}\,-L_{\text{GAN}}(\eta^{*},\tau)+\lambda_{w}L_{w}(\eta^{*},\tau)+\lambda_{p}L_{p}(\eta^{*},\tau)+\lambda_{\text{C}}L_{\text{C}}(\tau,\psi)+\lambda_{\text{DC}}L_{\text{DC}}(\tau)$ (20)
@@
In practice, we optimized the objectives for the generator (Eq. 19) and the discriminator (Eq. 20) with batch estimates and stochastic gradient descent. Moreover, we used separate optimizers for generator and discriminator optimization. Appendix 6.4 contains additional information about the generative model training.
@@ 4 Results
Figure 6: Disentanglement performance with encoder model for age-camera setup. We trained encoder models on retinal images, where the first subspace $w_{\text{age}}$ encoded the age class $y_{\text{age}}$, and the second subspace $w_{\text{cam}}$ encoded the camera class $y_{\text{cam}}$. In panels a and b, we present the improvement in kNN classifier accuracy over chance level accuracy. As a baseline model in a, we only trained linear classifiers on top of subspaces. In b, we further disentangle the subspaces using our distance correlation loss. Panel c illustrates the distance correlation measure between subspaces $w_{\text{age}}$ and $w_{\text{cam}}$ for both the baseline method and the method incorporating additional distance correlation minimization. In panels d and e, we provide t-SNE visualizations comparing a baseline model with a model incorporating disentanglement loss, respectively.
We conducted empirical evaluations to assess the efficacy of our disentanglement loss in two distinct scenarios: (1) a predictive task and (2) a generative task, both performed on retinal fundus images. Our primary objective in both cases was to disentangle the technical factor camera from patient attributes age and ethnicity.
For each setting, we conducted two experiments, configuring the model to learn specific subspaces. In the initial experiment, the first subspace $w_{\text{age}}\in\mathbb{R}^{4}$ encoded age information, while the second subspace $w_{\text{cam}}\in\mathbb{R}^{12}$ captured camera-related details. In the second experiment, we designed the first subspace $w_{\text{ethn}}\in\mathbb{R}^{8}$ to encode ethnicity information, while the second subspace focused on camera-related features $w_{\text{cam}}\in\mathbb{R}^{12}$. The dimensionality of each subspace was selected based on the dimensions of the labels we aimed to encode.
We filtered for images where labels for these classes were available and ended up with 75,708 images (24,254 patients) for the first and 63,210 images (19,950 patients) for the second experiment (see Sec. 3.1 for details). For more information about predictive model training, see Appendix 6.3.
@@ 4.1 Learning disentangled subspaces with an encoder model
We first evaluated the effect of the distance correlation loss on the simplest representation learning setup for images: an encoder neural network that mapped retinal images to a latent subspace representation (Fig. 3). To encourage disentangled subspaces, we studied the effect of minimizing the distance correlation between them (Sec. 3.3). To this end, we compared 4 runs of two model configurations on the test data: a baseline model, in which we trained only linear classifiers for subspace encoding, and a model that additionally minimized the distance correlation between subspaces.
@@
Our experiments suggest that the age and camera subspaces were heavily entangled in the baseline setting when disentanglement was not specifically enforced (Fig. 6a), such that camera information was present in the age subspace and vice versa. Entanglement was visible by relatively high classification accuracy across subspaces (Fig.6a, off-diagonals). For example, camera type could be decoded from the age subspace $w_{\text{age}}$ with a performance of $28\%$ above chance level, compared to $45\%$ in the dedicated camera subspace. When enforcing disentanglement by additionally minimizing the distance correlation between subspaces the cross-subspace classification performance (off-diagonals in Fig. 6b) dropped, while the within-subspace classification was preserved (diagonals). In addition, the distance correlation between subspaces on the test data dropped as a direct consequence of our disentanglement loss. (Fig. 6c) as a direct consequence of minimizing our disentanglement loss. We could also observed the disentanglement effect qualitatively in 2D t-SNE visualizations of the learned representations (Fig. 6d, e for a baseline and disentangled model, respectively). In both the baseline and the disentangled model, the age subspace exhibited structure w.r.t. age, and the camera subspace encoded camera type well (diagonals in Fig. 6d, e). In the baseline model, however, the age subspace also strongly encoded camera information (first row, second column in d).
@@
Figure 7: Disentanglement performance with encoder model for ethnicity-camera setup. We trained encoder models on retinal images, where the first subspace $w_{\text{ethn}}$ encoded the ethnicity class $y_{\text{ethn}}$, and the second subspace $w_{\text{cam}}$ encoded the camera class $y_{\text{cam}}$. In panels a and b, we present the improvement in kNN classifier accuracy over chance level accuracy. As a baseline model in a, we only trained linear classifiers on top of subspaces. In b, we further disentangle the subspaces using our distance correlation loss. Panel c illustrates the distance correlation measure between subspaces $w_{\text{age}}$ and $w_{\text{cam}}$ for both the baseline method and the method incorporating additional distance correlation minimization. In panels d and e, we provide t-SNE visualizations comparing a baseline model with a model incorporating disentanglement loss, respectively.
@@
In the second experiment, we observed an entanglement between the ethnicity and camera subspaces when employing the baseline method (Fig. 7a). For the baseline method, comparable attribute performance was evident across subspaces (columns Fig. 7a). For instance, camera decoding in the ethnicity subspace was at 44% accuracy improvement over chance level, compared to 47% for the dedicated camera subspace. Additionally, ethnicity decoding exhibited marginal improvement over chance level, with only 9% and 8%, respectively. This discrepancy could be attributed to the imbalance of our dataset, with 71% of patients being of Latin American descent (Fig. 2c). Upon incorporating our disentanglement loss into the training process, distance correlation dropped notably (Fig. 7c). Moreover, cross-subspace classification performance dropped (off-diagonals in Fig. 7a), while within-subspace classification performance remained high (diagonals in Fig. 7a). The impact of improved disentanglement was also evident in t-SNE visualizations. The baseline model exhibited camera clustering within the ethnicity subspace (Fig. 7d), a phenomenon not observed in the model incorporating our disentanglement loss (Fig. 7e).

### v2

@@ 3.3 Disentanglement loss
If the predictive model (Fig. 3) was only composed of an encoder and subspace classifiers, subspaces could correlate and share information. Therefore, we additionally penalized the presence of shared information with a disentanglement loss by minimizing the distance correlation (dCor) between unique subspace pairs:
@@
$L_{\text{DC}}(\theta)$ $=\frac{2}{K(K-1)}\sum_{i=1}^{K}\sum_{j=1}^{i-1}\text{dCor}(w_{i},w_{j}).$ (4)
Distance correlation (Székely et al., 2007) measures the dependence between random vectors of arbitrary dimension. It is analogous to Pearson’s correlation but can also measure nonlinear dependencies. Moreover, distance correlation is bounded in the $[0,1]$ range and is zero only if the random vectors are independent.
Examples of different dependence measures.
We provided a few 2D toy examples in Appendix A, where we compared the performance of minimizing distance correlation with minimizing linear dependence measures. When the relationship between the random variables was nonlinear, only minimizing distance correlation showed an effect.
We use the empirical distance correlation (Székely et al., 2007) to measure correlation between batch samples of subspace tensors:
$\text{dCor}(w_{1},w_{2})$ $=\frac{\text{dCov}(w_{1},w_{2})}{\sqrt{\text{dCov}(w_{1},w_{1})\,\text{dCov}(w_{2},w_{2})}}.$ (5)
We assumed n-dimensional batches of subspace vectors $w_{1}\in\mathbb{R}^{n\times d_{1}}$ and $w_{2}\in\mathbb{R}^{n\times d_{2}}$. Here, dCov is the distance covariance
$\text{dCov}^{2}(w_{1},w_{2})$ $=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}A_{i,j}B_{i,j}.$ (6)
Each matrix element $a_{i,j}$ of $A$ is the Euclidean distance between two samples
$A\in\mathbb{R}^{n\times n},\,a_{i,j}$ $=\|w_{1}^{(i)}-w_{1}^{(j)}\|_{2},\,i,j=1,2,\dots,n$ (7)
after subtracting the mean of row $i$ and column $j$, as well as the matrix mean. B is similarly calculated for $w_{2}$.
@@ 3.4 Generative image model
For the generative model, the overall goal was to generate realistic fundus images from a disentangled latent space. We considered the image generation task by learning the mapping between two domains $M:\mathcal{X}\rightarrow\mathcal{\hat{X}}$ with a latent space model. We split $M$ into two components, $M=d\circ g$, where the encoder/discriminator $d$ maps to an intermediate latent representation $d:\mathcal{X}\rightarrow\mathcal{W}$ and the generator $g:\mathcal{W}\rightarrow\mathcal{\hat{X}}$ maps to the output image domain.
@@
We chose to work with the StyleGAN2 architecture (Karras et al., 2020), a state-of-the-art generative adversarial network for high-fidelity and high-resolution images. The StyleGAN architecture has two features that makes it a potentially good candidate for disentanglement: (1) it consists of a mapping network $m:\mathcal{Z}\rightarrow\mathcal{W}$ that maps a latent vector into an intermediate space, which then controls the styles in each convolutional layer in the generator with adaptive instance normalization, and (2) it can control scale-specific image generation by a progressively growing generator.
@@
We trained the generative image model with the standard minimax GAN loss together with StyleGAN’s path length regularization $P_{L}$ (Karras et al., 2020) on the generator and an R1 regularizer (Mescheder et al., 2018) as a gradient penalty on the discriminator for real data:
$L_{\text{GAN}}(\eta,\tau)$ $=\,\mathbb{E}_{x\in\mathbb{X}}\left[\log(d_{\tau,0}(x))\right]+\mathbb{E}_{z\in\mathbb{Z}}\left[\log(1-d_{\tau,0}(g_{\eta}(w)))\right]+\lambda_{P_{L}}P_{L}(\eta)-\lambda_{R_{1}}R_{1}(\tau),$ (8)
$\eta^{*}$ $=\operatorname*{arg\,min}_{\eta}\,L_{\text{GAN}}(\eta,\tau^{*}),$ (9)
$\tau^{*}$ $=\operatorname*{arg\,min}_{\tau}\,-L_{\text{GAN}}(\eta^{*},\tau).$ (10)
@@
Here, $d_{\tau,0}(x)$ is the discriminator’s estimate of the probability that the real data instance $x$ is real ($0\leq d_{\tau,0}(\cdot)\leq 1$), $g_{\eta}(w)$ denotes the generator’s output given $w$, and $d_{\tau,0}(\hat{x})$ is the discriminator’s estimate of the probability that a fake instance $\hat{x}=g_{\eta}(w)$ is real (Fig. 4).
@@
Figure 4: Generative image model with subspace constraints. For each image, we sample three individual latent codes from a standard normal distribution $z_{k}\sim\mathcal{N}(0,I_{d_{k}})$ and map them to intermediate latent spaces. The concatenation of all three subspaces $w=[w_{1},w_{2},w_{3}]\in\mathbb{R}^{d_{1}+d_{2}+d_{3}}$ serves as input to the generator $g_{\eta}:\mathcal{W}\rightarrow\mathcal{X}$. For GAN-inversion, we extend our discriminator with two additional heads: one for latent representation $\hat{w}$ and another for a pixel feature vector $p$ for image encoding and pixel space reconstruction, respectively. For subspace learning, we introduce individual mapping networks for each subspace and incorporate label information for real images through classification heads. To prevent subspace correlation, we add a penalty to the classification tasks using the average distance correlation (dCor) between the subspaces $\hat{w}_{k}$.
@@
As GANs were not originally designed for representation learning, they do not provide an encoder model. Therefore, we extended the StyleGAN2 architecture to embed real images in the latest space for representation learning. Inspired by the work of invGAN (Ghosh et al., 2022) for GAN-inversion, we also trained the discriminator as an encoder. We added two independent, fully connected layers and returned an estimate of the latent representation $\hat{w}$ for our first inversion loss
$L_{w}(\eta,\tau)$ $=\mathbb{E}_{z\in\mathbb{Z}}\left[\|w-\hat{w}\|_{2}^{2}\right]=\mathbb{E}_{z\in\mathbb{Z}}\left[\|w-d_{\tau,1}(g_{\eta}(w))\|_{2}^{2}\right].$ (11)
@@
The main methodological contribution of this work is to extend the StyleGAN2 architecture (with GAN-inversion for representation learning) to an independent subspace learner. Therefore, we modified the architecture to sample individual latent codes from a standard normal distribution $z_{k}\sim\mathcal{N}(0,I_{d_{k}})$ and learned individual 8-layer mapping networks for each subspace $m_{\eta}^{k}:\mathcal{Z}_{k}\rightarrow\mathcal{W}_{k}$ (Fig. 4, left), where each subspace can have a different dimensionality $w_{k}\in\mathbb{R}^{d_{k}}$. In setting up the generative image model, we assumed that we do not know all the underlying factors of variation or have access to their labels. Therefore, we provided the model with a free patient-identity subspace $w_{3}$ (Fig. 4), in which the encoder/discriminator can capture further information required to reconstruct high-resolution retinal fundus images. As in the encoder model (Sec. 3.2), we encoded label information for real images by training linear classification heads $C_{\psi_{k}}$ for the first two subspaces (Fig. 4, right). To avoid subspace correlation, we penalized the classification task with the mean of all distance correlation measures between subspaces $\hat{w}_{k}$ that the discriminator outputs.
@@
$(\tau^{*},\psi^{*})$ $=\operatorname*{arg\,min}_{\tau,\psi}\,-L_{\text{GAN}}(\eta^{*},\tau)+\lambda_{w}L_{w}(\eta^{*},\tau)+\lambda_{p}L_{p}(\eta^{*},\tau)+\lambda_{\text{C}}L_{\text{CE}}(\tau,\psi)+\lambda_{\text{DC}}L_{\text{DC}}(\tau)$ (14)
@@
In practice, we optimized the objectives for the generator (Eq. 13) and the discriminator (Eq. 14) with batch estimates and stochastic gradient descent. Moreover, we used separate optimizers for generator and discriminator optimization. Appendix C contains additional information about the generative model training.
Figure 5: Encoder disentanglement performance for age-camera setup. We trained encoder models on retinal images, where the first subspace $w_{\text{age}}$ encoded age $y_{\text{age}}$, and the second subspace $w_{\text{cam}}$ encoded camera information $y_{\text{cam}}$. In panels a and b, we show the accuracy improvement over chance level accuracy. As a baseline model in a, we only trained linear classifiers on top of subspaces. In b, we further disentangle the subspaces using our distance correlation loss. Panel c shows the distance correlation measure between subspaces for both the baseline method and our proposed method for subspace disentanglement. In panels d and e, we provide t-SNE visualizations comparing a baseline model with a model with disentanglement loss, respectively.
@@ 4 Results
We conducted empirical evaluations to assess the efficacy of our disentanglement loss in two distinct scenarios: (1) a predictive task and (2) a generative task, both performed on retinal fundus images. Our primary objective in both cases was to disentangle camera as a technical factor from the patient attributes age and ethnicity.
To evaluate the success, we need to define a quantitative measure of disentanglement. While there is no universally accepted definition of disentanglement, most agree on two main points (Eastwood and Williams, 2018; Carbonneau et al., 2022): (1) factor independence and (2) information content. Factor independence means that a factor affects only a subset of the representation space, and only this factor affects this subspace. We refer to this property as modularity. As another criterion for factor independence some argue that the affected representation space by a factor should be as small as possible, a principle known as compactness. The second agreed-upon criterion is the information content or explicitness. In essence, a disentangled representation should completely describe all factors of interest. To assess disentanglement, three distinct families of metrics have been proposed: intervention-based, predictor-based, and information-based metrics (Carbonneau et al., 2022).
In our case, we opted for a predictor-based metric to evaluate both modularity and explicitness, excluding compactness because we fixed the subspace dimensions. For our predictor, we employed a k-nearest neighbors (kNN) classifier with k=30. Using this classifier, we generated a confusion matrix, comparing subspaces with associated labels, for example ethnicity and camera (Fig. 6). Given the presence of unbalanced classes, we report the accuracy improvement over chance level accuracy. For the confusion matrix, modularity is evident by a high classification accuracy improvement for one factor (e.g. ethnicity) in only one subspace. Explicitness, on the other hand, is evident when the accuracy improvement is high across the entire matrix. Consequently, when both modularity and explicitness are effectively addressed, the confusion matrix has high values along the main diagonal as well as low values on the off-diagonals (Fig. 6).
For both the predictive and the generative task, we conducted two experiments, configuring the model to learn specific subspaces. In the initial experiment, the first subspace $w_{\text{age}}\in\mathbb{R}^{4}$ encoded age information, while the second subspace $w_{\text{cam}}\in\mathbb{R}^{12}$ captured camera-related details. In the second experiment, we designed the first subspace $w_{\text{ethn}}\in\mathbb{R}^{8}$ to encode ethnicity information, while the second subspace focused again on camera-related features $w_{\text{cam}}\in\mathbb{R}^{12}$. The dimensionality of each subspace was selected based on the dimensions of the labels we aimed to encode.
We filtered for images where labels for these classes were available and ended up with 75,708 images (24,254 patients) for the first and 63,210 images (19,950 patients) for the second experiment (see Sec. 3.1 for details).
@@ 4.1 Learning disentangled subspaces with an encoder model
We first evaluated the effect of the distance correlation loss on the simplest representation learning setup for images: an encoder neural network that mapped retinal images to a latent space (Fig. 3). To encourage disentangled subspaces, we studied the effect of minimizing the distance correlation between them (Sec. 3.3). To this end, we compared 4 runs of two model configurations on the test data: a baseline model, for which we trained only linear classifiers for subspace encoding, and a model that additionally minimized the distance correlation between subspaces. For more information about the predictive model training, see Appendix B.
@@
Our experiments showed that the age and camera subspaces were heavily entangled in the baseline setting when disentanglement was not specifically enforced (Fig. 5a), such that camera information was present in the age subspace and vice versa. Entanglement was visible by relatively high classification accuracy across subspaces (Fig.5a, off-diagonals). For example, camera type could be decoded from the age subspace $w_{\text{age}}$ with a performance of $28\%$ above chance level, compared to $45\%$ in the camera subspace. When enforcing disentanglement by additionally minimizing the distance correlation between subspaces the cross-subspace classification performance (off-diagonals in Fig. 5b) dropped, while the within-subspace classification was preserved (diagonals). In addition, the distance correlation between subspaces on the test data dropped as a direct consequence of our disentanglement loss (Fig. 5c). We could also observe the disentanglement effect qualitatively in 2D t-SNE visualizations of the learned representations. In both the baseline and the disentangled model, the age subspace exhibited structure w.r.t. age, and the camera subspace encoded camera type well (diagonals in Fig. 5d, e). In the baseline model, however, the age subspace also strongly encoded camera information (first row, second column in d).
@@
Figure 7: Encoder disentanglement performance for ethnicity-camera setup. We trained encoder models on retinal images, where the first subspace $w_{\text{ethn}}$ encoded ethnicity $y_{\text{ethn}}$, and the second subspace $w_{\text{cam}}$ encoded camera information $y_{\text{cam}}$. In panels a and b, we show accuracy improvement over chance level accuracy. As a baseline model in a, we only trained linear classifiers on top of subspaces. In b, we further disentangle the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces of our two model configurations. In panels d and e, we provide t-SNE visualizations comparing a baseline model with a model with disentanglement loss, respectively.
@@
In the second experiment, we observed an entanglement between the ethnicity and camera subspaces for the baseline method (Fig. 7a). For the baseline method we observed comparable attribute performance across subspaces (columns Fig. 7a). For example, camera decoding in the ethnicity subspace showed an accuracy improvement of 44% over chance level, compared to 47% for the dedicated camera subspace. In addition, ethnicity decoding only showed a marginal improvement over chance level, with 9% and 8%, respectively. This discrepancy may be due to the imbalance of our dataset, with 71% of the patients being Latin American (Fig. 2c). After incorporating our disentanglement loss, distance correlation dropped notably (Fig. 7c). Moreover, cross-subspace classification performance decreased (off-diagonals in Fig. 7a), while within-subspace classification performance remained high (diagonals in Fig. 7a). The impact of the improved disentanglement was also evident in the t-SNE visualizations. The baseline model exhibited camera clustering within the ethnicity subspace (Fig. 7d), a phenomenon not observed in the model with our disentanglement loss (Fig. 7e).